Show that P^-1HP = 3à —3 matrix



Use cofactor method to find the inverse of P


1
Expert's answer
2022-01-25T04:43:00-0500

We shall find the inverse of matrix "p," denoted as "p^{-1}," using the cofactor method

Say the "3" X "3" matrix is given by

"p=\begin{bmatrix}
 2 & 1 & 0 \\
 1 & -1 & 1 \\
 3 & 2 & 1
\end{bmatrix}"


Next, we find the determinant of "p" denoted as "\mid p\mid"


"\mid p \mid=2\begin{vmatrix}
 -1 & 1 \\
 2 & 1
\end{vmatrix}-1\begin{vmatrix}
 1 & 1 \\
 3 & 1
\end{vmatrix}+0\begin{vmatrix}
 1 & -1 \\
 3 & 2
\end{vmatrix}"


"=2(-3)-1(-2)+0"


"= -6+2"


"=-4"

Thus, we have "\mid p \mid= -4"


Now, We shall find the cofactor which is given by

"= \begin{bmatrix}
+\begin{vmatrix}
 -1 & 1 \\
 2 & 1
\end{vmatrix} & 
-\begin{vmatrix}
 1 & 1 \\
 3 & 1
\end{vmatrix} & 
+ \begin{vmatrix}
 1 & -1 \\
 3 & 2
\end{vmatrix}\\

- \begin{vmatrix}
 1 & 0 \\
 2 & 1
\end{vmatrix} & +\begin{vmatrix}
 2 & 0 \\
 3 & 1
\end{vmatrix} & -\begin{vmatrix}
 2 & 1 \\
 3 & 2
\end{vmatrix}\\
+ \begin{vmatrix}
 1 & 0 \\
 -1 & 1
\end{vmatrix} & -\begin{vmatrix}
 2 & 0 \\
 1 & 1
\end{vmatrix} & +\begin{vmatrix}
 2 & 1 \\
 1 & 1
\end{vmatrix}
\end{bmatrix}"


"=\begin{bmatrix}
 -3 & 2 & 5 \\
 -1 & 2 & -1\\
 1 & -2 & -3
\end{bmatrix}"


Now, by finding the transpose of the above matrix and dividing by the determinant

we have,


"\displaystyle=\frac{1}{-4}\begin{bmatrix}
 -3 & -1 & 1 \\
 2 & 2 & -2\\
5 & -1 & -3
\end{bmatrix}=p^{-1}"


Hence, we have

"\displaystyle p^{-1}= \begin{bmatrix}
 \frac{3}{4} & \frac{1}{4} & \frac{-1}{4} \\
 \frac{-1}{2}& \frac{-1}{2} & \frac{1}{2}\\
 \frac{-5}{4} & \frac{1}{4} & \frac{3}{4}
\end{bmatrix}"

which is the inverse of the matrix by the cofactor method.